# Combinatorial metaplexes and centrality indices for identifying higher-order interactions

### v1

@@ 4.1  Background: centrality in weighted graphs
In this succeeding section, we extend the fully weighted graph centrality model to the CM by replacing scalar vertex weights with simplicial concentrations and edge weights with facet-induced adjacency strengths.
@@ 4.2  Adjacency and degree in the CM
In this section, we introduce a family of centrality maps defined on the set of $q$-simplices. Each centrality map assigns to a simplex a real value determined jointly by its incidence relations and the concentration map, and satisfies structural properties compatible with the CM framework. These maps induce a ranking on simplices and provide a quantitative characterisation of their structural roles within the CM, thereby distinguishing the CM from purely topological simplicial complexes.
We focus on centrality indices derived from the incidence structure (degree) and from the shortest-path structure (closeness and harmonic). In the CM framework, these indices are defined with respect to the weighted adjacency relations induced by the concentration map on $\mathcal{K}_{c}$. We now formalise this construction.
@@ 4.3  Weighted walks and distances
This definition reflects the modelling assumption that larger weights of mediating facets and destination simplices indicate stronger combinatorial influence in the formation of higher-order structure. Accordingly, the contribution of each step in the walk is taken to be inversely proportional to the product $a(\gamma_{l})\,a(\sigma_{i_{l+1}})$. Under this choice, steps involving larger weights contribute smaller increments to the total length.
@@
Let $\alpha\geq 0$. The combined shortest path distance between $\sigma_{i},\sigma_{j}\in S_{q}(\mathcal{K}_{c})$ is defined by
@@ 4.4  Weighted closeness centrality
Let $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The weighted closeness centrality of $\sigma_{i}$ is defined by
@@
Let $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The harmonic weighted closeness centrality of $\sigma_{i}$ is defined by
@@
Let $\alpha\geq 0$ and $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The combined weighted closeness centrality of $\sigma_{i}$ is defined by
@@
Analogously, the harmonic combined centrality is defined by
@@ 5  Conclusion
In this work, we introduce the combinatorial metaplex as a mathematical framework for modelling higher-order systems in which interaction structure and intrinsic unit properties are intrinsically coupled. In contrast to classical simplicial-complex-based models, where higher-order simplices arise solely from inclusion relations, the combinatorial metaplex integrates vertex-level intrinsic properties directly into the realisation of higher-dimensional interactions. The resulting structure preserves the combinatorial rigour of simplicial modelling while incorporating discrete intrinsic attributes of the constituent units.
A central contribution of the paper is the formal construction of the concentration map. Beginning with a vertex-level concentration assignment, we provide a constructive extension to all simplices that is compatible with the simplicial incidence structure. In particular, we prove that the total concentration of vertices equals the total concentration of facets, establishing a hierarchical conservation principle. This construction clarifies how vertex-level quantities propagate across dimensions and determine the strength and admissibility of higher-order interactions.
Within this framework, higher-order simplices are realised only when their aggregated boundary concentration exceeds a prescribed threshold. This criterion enables the identification of concentration-supported higher-order interactions during reconstruction, distinguishing them from simplices that would arise solely through combinatorial closure, as in clique-based constructions.
Building on the induced concentration structure, we define facet-mediated adjacency for non-facet simplices and introduce corresponding degree, closeness, and harmonic centrality indices. These measures respect the simplicial organisation while incorporating concentration-driven coupling. The associated one-parameter family of adjacency relations interpolates between purely structural connectivity and concentration-induced interaction.
The combinatorial metaplex therefore provides a unified setting in which interaction topology and intrinsic vertex properties jointly determine higher-order organisation. It may be viewed as a structurally constrained subclass of weighted simplicial complexes characterised by vertex-induced propagation and hierarchical conservation across dimensions. This constraint distinguishes it from arbitrary simplex-level weight assignments and ensures coherence between local intrinsic properties and global higher-order structure.
The framework has been developed in a purely theoretical setting. Extending it to empirical datasets and application domains constitutes a natural direction for future research.

### v2

@@ 4.1  Background: centrality in weighted graphs
In the succeeding section, we extend the fully weighted graph centrality model to the CM by replacing scalar vertex weight with simplex weight and edge weight with facet-induced adjacency strength.
@@ 4.2  Adjacency and degree in the CM
In this section, we introduce a family of centrality maps defined on the set of $q$-simplices. Each centrality map assigns to a simplex a real value determined jointly by its incidence structure and the weight map, and satisfies structural properties compatible with the CM framework. These maps induce a ranking on simplices and provide a quantitative characterisation of their structural importance within the CM, thereby distinguishing the CM from purely topological simplicial complexes.
We focus on centrality indices derived from the incidence structure (degree) and from the shortest-path distance (closeness and harmonic). In the CM framework, these indices are defined with respect to the weighted adjacency relations induced by the concentration map on $\mathcal{K}_{c}$. We now formalise this construction.
@@ 4.3  Weighted walks and distances
This definition reflects the modelling assumption that higher weights of mediating facets and destination simplices indicate stronger combinatorial influence in the formation of higher-order structure. Accordingly, the contribution of each step in the walk is taken to be inversely proportional to the product $a(\gamma_{l})\,a(\sigma_{i_{l+1}})$. Under this choice, steps involving larger weights contribute smaller increments to the total length.
@@
Let $\alpha\in[0,1]$. The combined shortest path distance between $\sigma_{i},\sigma_{j}\in S_{q}(\mathcal{K}_{c})$ is defined by
@@ 4.4  Weighted closeness centrality
Let $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The weighted closeness centrality (CC) of $\sigma_{i}$ is defined by
@@
Let $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The weighted harmonic centrality (HC) of $\sigma_{i}$ is defined by
@@
Let $\alpha\in[0,1]$ and $\sigma_{i}\in S_{q}(\mathcal{K}_{c})$. The combined weighted closeness centrality of $\sigma_{i}$ is defined by
@@
Analogously, the combined harmonic centrality is defined by
@@
5  Illustration and results
In this section, we illustrate the CM framework developed in Section 3 using a real-world network. Specifically, the CM is constructed from pairwise data obtained from the food web dataset available in the cheddar R package [29]. We consider the food web of Tuesday Lake, Michigan, United States of America, whose pairwise network consists of 50 vertices and 263 edges.
In this network, the discrete structure assigned to each vertex represents the collection of all individuals belonging to the corresponding species, while the weight denotes the total biomass of that species, measured in units of $mg\,m^{-3}$. Based on these vertex weights, we define the fractional contribution map, under which the weight of a (q1)(q-1)-simplex is distributed equally among all its incident $q$-simplices for $1\leq q\leq 3$. Using the threshold criterion introduced earlier, we identify the HoIs included in the CM, which we refer to as true HoIs in this work.
The clique complex of this network contains 50 0-simplices, 263 $1$-simplices, 170 $2$-simplices, 113 $3$-simplices, 49 $4$-simplices, 13 $5$-simplices and 2 $6$-simplices. Table 1 presents a comparison between the number of HoIs obtained from the clique complex and true HoIs obtained from the CM.
Dimension of a simplex No. of simplices in the clique complex No. of simplices in the CM
0-simplices 50 50
1-simplices 263 263
2-simplices 170 89
3-simplices 113 5
4-simplices 49 0
5-simplices 13 0
6-simplices 2 0
Total 610 407
Table 1: Comparison of the number of simplices across dimensions in the clique complex and the CM, highlighting that the CM identifies true HoIs.
(a) Clique complex
(b) CM
Figure 1: Comparison of the 2-skeletons of the clique complex and the CM. The boundaries of triangles in the clique complex are shown in blue, while the boundaries of true triangles in the CM are shown in red.
Dimension of a simplex No. of facets Sum of facet weights No. of non-facets Sum of non-facet weights
0-simplices 0 0 50 1171.372469
1-simplices 162 412.376647 101 758.9958
2-simplices 72 576.776620 17 158.8958
3-simplices 5 182.219202 0 0
Table 2: Summary of facet and non-facet simplices counts and their total weights across dimensions, verifying Theorem 3.5 by showing that the sum of vertex weights equals the sum of facet weights.
Edge Edge weight Simplicial degree: $k$ ($\alpha=0$) Combined degree: $D^{\alpha}$ ($\alpha=0.5$) Weighted degree: $D^{1}$ ($\alpha=1$)
(47,48) 34.8970 21 466.2371 10351.2867
(48,50) 21.4450 21 418.8909 8355.6944
(47,50) 16.5480 21 403.2099 7741.8187
(41,47) 17.6106 40 506.3649 6410.1363
(28,47) 16.6017 9 216.0144 5184.6920
(46,47) 16.6851 39 441.1313 4989.6619
(28,50) 3.1497 5 156.9601 4927.2966
(31,47) 15.1792 9 203.2567 4590.3667
(32,47) 15.3263 15 261.3505 4553.6043
(31,50) 1.7272 5 148.6004 4416.4158
$\vdots$
(40,49) 1.6479 2 19.2811 185.8814
(27,30) 1.0547 2 18.8169 177.0369
(30,49) 2.3517 2 18.4353 169.9292
(30,38) 0.8344 2 18.2916 167.2912
(26,30) 0.8132 2 18.2405 166.3572
Table 3: Top 10 and bottom 5 edges ranked by weighted degree $D^{1}$ ($\alpha=1$). The table presents edge weights together with the simplicial degree $k$ ($\alpha=0$) and the combined degree $D^{\alpha}$ ($\alpha=0.5$).
Edge Edge weight $\mathbf{CC^{0}}$ $\mathbf{CC^{0.5}}$ $\mathbf{CC^{1}}$
$(41,47)$ 17.6106 0.006061 0.042212 0.160132
$(47,48)$ 34.8970 0.005587 0.041699 0.159768
$(46,47)$ 16.6851 0.006024 0.041559 0.159610
$(41,48)$ 22.5076 0.004717 0.039746 0.158834
$(48,50)$ 21.4450 0.004292 0.039209 0.158776
$(41,46)$ 4.2957 0.004405 0.039274 0.158700
$(46,48)$ 21.5821 0.004630 0.039435 0.158689
$(47,50)$ 16.5480 0.005587 0.039007 0.158659
$(28,47)$ 16.6017 0.004000 0.038666 0.158603
$(31,50)$ 1.7272 0.003802 0.038249 0.158565
$\vdots$
$(40,49)$ 1.6479 0.003636 0.028457 0.137050
$(27,30)$ 1.0547 0.003846 0.028028 0.136933
$(26,30)$ 0.8132 0.003846 0.027803 0.136921
$(30,38)$ 0.8344 0.003846 0.027822 0.136899
$(30,49)$ 2.3517 0.003968 0.027978 0.134071
Table 4: Top 10 and bottom 5 edges ranked according to closeness centrality ($CC^{1}$). The corresponding values for $\alpha=0,0.5,1$, together with the edge weights, are also reported.
Edge Edge weight $\mathbf{HC^{0}}$ $\mathbf{HC^{0.5}}$ $\mathbf{HC^{1}}$
$(47,48)$ 34.8970 60.5000 795.0753 13893.3599
$(47,50)$ 16.5480 60.5000 719.8201 13658.1812
$(48,50)$ 21.4450 51.5000 728.4055 13610.8367
$(28,50)$ 3.1497 41.1667 680.7542 13354.9942
$(28,47)$ 16.6017 44.6667 688.3476 13024.0892
$(31,50)$ 1.7272 41.1667 669.2698 12828.6989
$(28,48)$ 21.4987 41.1667 661.6855 12540.4101
$(31,47)$ 15.1792 44.6667 673.5648 12501.1194
$(32,50)$ 1.8743 42.1667 662.8607 12417.5455
$(32,47)$ 15.3263 50.1667 675.0535 12130.6257
$\vdots$
$(40,49)$ 1.6479 38.3333 354.9303 3278.5304
$(27,30)$ 1.0547 41.0000 336.8375 3044.8693
$(30,38)$ 0.8344 41.0000 332.8572 2966.6504
$(26,30)$ 0.8132 41.0000 332.4688 2958.9547
$(30,49)$ 2.3517 42.1667 335.1990 2921.9739
Table 5: Top 10 and bottom 5 edges ranked according to harmonic centrality ($HC^{1}$). The corresponding values for $\alpha=0,0.5,1$, together with the edge weights, are also reported.
Figure 1 illustrates the 2-skeletons of the clique complex and the CM for comparison: blue lines indicate the boundaries of triangles present in the clique complex, whereas red lines indicate the boundaries of true triangles present in the CM.
Table 2 verifies Theorem 3.5 by reporting the number and total weights of facets and non-facet simplices across dimensions. The results confirm that the total weight of the vertices equals the total weight of the facets.
The analysis of the data using degree distributions of edges is shown in Table 3. It shows the top 10 edges having the highest weighted degree, along with their weights, simplicial degree, and combined degree ($\alpha=0.5$).
Table 4 presents the top ten and bottom five edges with respect to closeness centrality under the weighted walk for $\alpha=0$, $0.5$, and $1$. The edges are ranked according to the values obtained for $\alpha=1$, and their corresponding weights are also reported. The shortest-path distances are computed using Dijkstra’s algorithm.
Table 5 presents the top ten and bottom five edges with respect to harmonic centrality across the same range of $\alpha$ values. The edges are ranked according to the values obtained for $\alpha=1$, and their corresponding weights are also reported.
We observe from Tables 3, 4, and 5 that among the top 10 edges, four edges are common across all three centrality indices, while the remaining edges vary. However, the last five edges are the same in all three centralities, with only a change in their order.
The following observations could be made from the above discussions:
  1. (i)
    A comparison between the simplices arising from the clique complex and those constructed under the CM shows that the CM yields a more precise and structurally meaningful representation for the inference of higher-order interactions (HoIs).
  2. (ii)
    The weight distribution across dimensions verifies the conservation principle established in the theorem, namely that the total weight assigned to vertices coincides exactly with the total weight carried by the facets.
  3. (iii)
    The centrality measures computed for different values of $\alpha$ illustrate how the relative importance of simplices shifts as the model transitions from a purely combinatorial regime to a weight-dominated regime.
@@ 6  Conclusion
In this work, we introduce the combinatorial metaplex as a mathematical framework for modelling higher-order networks, in which interaction structure and intrinsic unit properties are coupled. In contrast to classical clique complex-based models, where HoIs included solely from inclusion relations, the combinatorial metaplex couples vertex-level intrinsic properties directly into the inclusion of HoIs.
A central contribution of the paper is the formal construction of the concentration map presented in Section 3. Starting from a vertex concentration assignment map $a:V\to\mathbb{Q}^{+}$, we extend this assignment to a weight map on the simplicial complex. The extension is defined via fractional contribution maps that distribute the weight of lower-dimensional simplices among the higher-dimensional simplices that contain them. The resulting procedure defines a coherent propagation of intrinsic quantities across the simplicial hierarchy. A key structural result is the conservation relation established in Theorem 3.5, which shows that the total weight assigned to the vertices equals the total induced weight of the facets. This result establishes a vertex–facet conservation principle, whereby the total vertex weight is exactly preserved in the induced facet weights.
Within this framework, higher-dimensional simplices are included only when the total weight of their boundary simplices exceeds a prescribed threshold. This rule yields a deterministic procedure for constructing simplicial complexes from weighted vertex data. Unlike clique-based constructions, which include every combinatorially closed simplex, the CM includes only those supported by sufficient boundary weight. This distinction separates the true HoIs captured by the CM from the HoIs that arise solely through combinatorial closure in the clique complex.
Using the induced weights on simplices and the facet-mediated adjacency between simplices, we develop corresponding centrality indices for the $q$-simplices, $0\leq q\leq\dim(\Delta)-1$, in Section 4. In particular, we introduce simplicial and weighted degree centralities, together with a one-parameter family $D^{\alpha}$ that interpolates between the simplicial and weighted degrees. Weighted walks and shortest-path distances between simplices then give rise to the closeness and harmonic centrality indices $CC^{\alpha}$ and $HC^{\alpha}$, which quantify the importance of simplices in the CM.
To illustrate the proposed framework, we construct a CM from the Tuesday Lake food-web dataset using vertex-level biomass values, as described in Section 5. This example demonstrates how the threshold rule determines the inclusion of true HoIs and how the concentration map $a$ subsequently induces weight on these simplices. A comparison with the corresponding clique complex shows that the CM identifies a reduced set of HoIs supported by sufficient boundary weight, thereby filtering out simplices that arise solely from combinatorial closure. This highlights the role of intrinsic vertex properties in governing the inclusion and structure of higher-order networks.
The evaluation of the proposed degree, closeness, and harmonic centrality indices in the illustrative example demonstrates how the weight-driven simplicial structure influences the relative importance of simplices in the CM. In particular, the results show that the tuning parameter regulates the balance between the simplicial topology and the vertex weights in determining centrality values. As the parameter varies, the centralities transition between behaviour governed primarily by the simplicial structure and behaviour influenced more strongly by the vertex weights. Together, these observations illustrate how the combinatorial metaplex provides a principled framework for constructing and analysing higher-order networks from weighted vertex data.
Declaration of competing interest
The authors have no conflicts of interest to declare that are relevant to the content of this article.
AI-Assistance Disclosure
The authors used an artificial intelligence language model (ChatGPT, developed by OpenAI) for limited assistance in rephrasing, language polishing, and basic grammatical correction of certain sentences. In addition, ChatGPT was used to assist in generating and refining computational code for implementing illustrative examples of the proposed framework using Google Colab. All mathematical results, algorithms, theoretical developments, interpretations, and scientific conclusions presented in this manuscript were entirely authored, verified, and approved by the authors.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.